OpenStudy (anonymous):

Simplify. Express the product as a radical expression. ^3sqrtx^2 times ^4sqrtx

OpenStudy (anonymous):

\[\large \sqrt[3]{x^2}\times \sqrt[4]{x}\]??

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

start with \[\large x^{\frac{2}{3}}\times x^{\frac{1}{4}}\] then add the exponents

OpenStudy (anonymous):

so x^2 3/7 ?@satellite73

OpenStudy (anonymous):

oh hello no!!

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}\neq \frac{a+c}{b+d}\]you got to add the fractions correctly even if they are in the exponent

OpenStudy (anonymous):

): math is not my stongsuit

OpenStudy (anonymous):

btw i means "f" no but i cannot say that to quote your bio "flutter no"

OpenStudy (anonymous):

flutter ahahah

OpenStudy (anonymous):

\[\frac{a}{b}+\frac{c}{d}=\frac{ad+bc}{bd}\] that is how you add

OpenStudy (anonymous):

so \[\frac{2}{3}+\frac{1}{4}=?\]

OpenStudy (anonymous):

you got this? let me know what you get adding fractions is something math teachers love to make you do also subtracting them

OpenStudy (anonymous):

11/12

OpenStudy (anonymous):

bingo

OpenStudy (anonymous):

don't never forget

OpenStudy (anonymous):

would the answer be ^12sqrtx^11

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

\[\sqrt[12]{x^{11}}\] yes

OpenStudy (anonymous):

you got this?

OpenStudy (anonymous):

yes thank you @satellite73

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

now flutter off and do something more productive on friday night than math